OpenStudy (anonymous):

Hello everyone! Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4.

OpenStudy (anonymous):

y=mx + b

OpenStudy (anonymous):

y - 2 = 4(x + 6) y + 6 = 4(x - 2) y + 2 = 4(x - 6) y - 6 = 4(x + 2)

OpenStudy (anonymous):

i think its the first one..

OpenStudy (anonymous):

I was thinking that too. Thanks

OpenStudy (anonymous):

@Tulyes your welcome.